A richshaw driver moving with a speed of 36 km/ h,notices a child standing in the middle of the road and brings his rickshaw to rest in 4 sec.just in time to save the child.What is the average retarding force on the rickshaw? The mass of the rickshaw is 400 kg and the mass of the driver is 65 kg.
1
Expert's answer
2017-04-04T15:44:07-0400
Initial speed of the three-wheeler is u=36 km/h = 10m/s; final speed v = 0 m/s.
Total mass of the system: M = 400 + 65 = 465 kg.
According to the first law of motion: v = u + at
a = (v-u)/t = (0-10)/4 = -2.5 m/s2
F = Ma = 465 ·(-2.5) = -1162.5 N
The negative sign indicates that this force acts against the direction of motion of the three-wheeler.